OpenStudy (anonymous):

\[X: \sqrt{3x}+3=5-x\]

jhonyy9 (jhonyy9):

you need solving for x , yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

in first step add to both sides -3

OpenStudy (anonymous):

i am going to make a guess that it is \[\sqrt{x+3}=5-x\]

OpenStudy (anonymous):

no its the way i posted it

jhonyy9 (jhonyy9):

sqrt3x +3 -3 =5-3-x sqrt3x =2-x so now squaring the both sides what will get ?

OpenStudy (anonymous):

then i will be quiet, jhonyy has it

jhonyy9 (jhonyy9):

so than 3x=(2-x)(2-x) can you continue ?

OpenStudy (anonymous):

no I'am totally lost

jhonyy9 (jhonyy9):

where and why ?

OpenStudy (anonymous):

I'm going to fail this class I'm not good in math at all can you have 2 answers

OpenStudy (anonymous):

I end up w 2 or 11

jhonyy9 (jhonyy9):

so 3x=x^2 -4x +4 x^2 -7x +4=0 x_1,2 =(7+/- sqrt(49-16))/2 =(7+/-sqrt33)/2

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

No I guess Im not going to get it my choices are 2, 11, or 2 and 11 or no solution

OpenStudy (anonymous):

It is multiple choice

jhonyy9 (jhonyy9):

so if you will check it with these roots will get those right ?

jhonyy9 (jhonyy9):

for 2 you will get 6=0

OpenStudy (anonymous):

well its a b c or d it is multiple choice so then there is no real solution

OpenStudy (anonymous):

given the are A of a circle the formula r= \[\sqrt{a/3.14}\] can be used to calculate the radius of the circle. Fine the area when the radius is 18 8.2 cm

jhonyy9 (jhonyy9):

area of a circle = pir^2 so from this result that r=sqrt(A/pi) so what you nhave wrote there yes is right sure